OpenStudy (anonymous):

If 0° < θ < 90° and cosθ= (5√2)/8, then tanθ =?

OpenStudy (anonymous):

well all u gotta do is find what theta equals and then plug that in to tan (theta)

OpenStudy (anonymous):

\[\cos^{-1} 5\sqrt{2}\div8 =\theta\]

OpenStudy (anonymous):

\[\theta = 27.88 \] which is approximately 28

OpenStudy (anonymous):

tan(28)=.53